laundry processed 120 kg & is expected to grow to 132 kg. This growth will continue at the same percentage rate for next 7years . Currently hospital , is considering two options, the purchase of machine A or the rental of machine B. given information: Machine A – purchase Annual capacity £180 Material cost per kilogram £2 Labour cost per kilogram £3 Fixed costs per annum £20 Life of machine 3 years Capital cost £60 Depreciation per annum £20 Machine B – rent Annual capacity (kilograms) £170 Material cost per kilogram £1.8 Labour cost per kilogram £3.4 Fixed costs per annum £18 Rental per annum £20 Rental agreement 3 years Depreciation per annum nil 1. The hospital is able to call on an outside laundry if there is either a breakdown. The charge would be £10 per kilogram of washing. 3. Machine A will have no residual value at any time. 4. Present machine can be sold £10 cash 6. discount rate is 15%. (a) evaluate the two options for operating the laundry, using discounted cash flow techniques;
1
Expert's answer
2019-03-21T12:52:48-0400


Processing capacity per year

y1=120

y2=132

y3= 132*(132/120)=145.2

y4=159.7

y5=175.7

y6=193.3

y7=212.6


option A

Annual costs

AC1=120*(2+3)+20-20=600

AC2=132*(2+3)+20-20=660

AC3=145.2*(2+3)+20-20+60=786


option B

Annual costs

AC1=120*(1.8+3.4)+18+20=662

AC2=132*(1.8+3.4)+18+20=724.4

AC3=145.2*(1.8+3.4)+18+20=793.04


Discounted cash flow

A=180+600*(1/(1.15^1))+660*(1/(1.15^2))+786*(1/(1.15^3))=1717.60

B=662*(1/(1.15^1))+724.4*(1/(1.15^2))+793.04*(1/(1.15^3))=1644.84



As a result of the calculation of discounted cash flows, will be better option B, since the cost of the laundry will be less than in option A.